OpenStudy (anonymous):

Simplify. 4^√2(√2)^2

OpenStudy (anonymous):



zepdrix (zepdrix):

Are you familiar with rewriting stuff with rational exponents? Example:\[\Large\rm \sqrt{x}=x^{1/2}\]

OpenStudy (anonymous):

umm..no..:(

zepdrix (zepdrix):

\[\Large\rm \sqrt[4]{2}\left(\color{orangered}{\sqrt{2}}\right)^2\]We could apply that here:\[\Large\rm \sqrt[4]{2}\left(\color{orangered}{2^{1/2}}\right)^2\]And here:\[\Large\rm \color{orangered}{\sqrt[4]{2}}\left(2^{1/2}\right)^2=\color{orangered}{2^{1/4}}\left(2^{1/2}\right)^2\]

zepdrix (zepdrix):

And apply some exponent rules to simplify.